# Open Packing in Graphs: Bounds and Complexity

### v1

@@ Appendix A Open Packing
$V(G_{r})=(\cup_{i=1}^{r}\{x_{i},y_{i},z_{i}\})\cup\{u,v\}$ and $E(G_{r})=(\cup_{i=1}^{r}\{x_{i}y_{i},y_{i}z_{i},z_{i}u\})\cup\{uv\}$. On the contrary assume that there exists a $(P_{4}\cup rK_{1})$ say $H$ in $G_{r}$. Let $\{a,b,c,d\}\in V(H)$ such that $G_{r}[{a,b,c,d}]\cong P_{4}$. Note that every $P_{4}$ in $G_{r}$ consists of the vertex $u$. So, $u\in\{a,b,c,d\}$. Similarly, observe that every $P_{4}$ in $G$ consists of exactly one vertex labelled $y_{i}$ for some $i=1,2\ldots,r$. So, $y_{i}\in\{a,b,c,d\}$ for some $i=1,2,\ldots,r$. W.L.O.G., assume that $i=1$. Then, $\{a,b,c,d\}\in\{\{v,u,z_{1},y_{1}\},\{z_{j},u,z_{1},y_{1}\},\{u,z_{1},y_{1},x_{1}\}\}$. Since $H\cong(P_{4}\cup rK_{1})$ and $\{a,b,c,d\}\subseteq V(H)$ induces the $P_{4}$, $G_{r}[V(H)\setminus\{a,b,c,d\}]\cong rK_{1}$ and no vertex in $(V(H)\setminus\{a,b,c,d\})$ is adjacent to a vertex in $\{a,b,c,d\}$. So, for every possibility of $\{a,b,c,d\}$, $(V(H)\setminus\{a,b,c,d\})\subseteq\{x_{2},y_{2},\ldots,x_{r},y_{r}\}$. But the independence number of $G_{r}[\{x_{2},y_{2},\ldots,x_{r},y_{r}\}]$ is $(r-1)$, a contradiction to $(V(H)\setminus\{a,b,c,d\})\subseteq\{x_{2},y_{2},\ldots,x_{r},y_{r}\}$ and $(V(H)\setminus\{a,b,c,d\})\cong rK_{1}$, i.e., the independence number of an induced subgraph of $G_{r}[\{x_{2},y_{2},\ldots,x_{r},y_{r}\}]$ is at least $r$.
@@
We use induction on $r$ to prove this lemma as well. We begin with the proof that $\rho^{o}(G)\leq 2$ for every connected $(P_{3}\cup K_{1})$-free graph $G$. Note that by Lemma 9, $\rho^{o}(G)\leq\gamma_{t}(G)\leq 3$, for every connected $(P_{3}\cup K_{1})$-free graph $G$. So, to prove the claim, it is enough to disprove the case $\rho^{o}(G)=3$. On the contrary, assume that there exists a $(P_{3}\cup K_{1})$-free graph $G$ with $\rho^{o}(G)=3$. Since the open packing number of a connected $P_{3}$-free graph is at most two by Remark 4 and $\rho^{o}(G)=3$, $G$ must contain $P_{3}$ as an induced subgraph. Let $D=\{x_{1},x_{2},x_{3}\}$ be a subset of $V(G)$ that induces a $P_{3}$ in $G$ with $x_{1}x_{2},x_{2}x_{3}\in E(G)$. Then, by the arguments in the proof of Lemma 9, $D$ is a total dominating set in $G$, i.e., $V(G)=\bigcup\limits_{i=1}^{3}N(x_{i})$. Then, for every open packing $S$ in $G$, $S=\bigcup\limits_{i=1}^{3}(S\cap N(x_{i}))$. We know that $|S\cap N(x_{i})|\leq 1$ for $i=1,2,3$. This together with the fact that $S=\bigcup\limits_{i=1}^{3}(S\cap N(x_{i}))$ implies that for an open packing $S$ in $G$ of cardinality three, $|S\cap N(x_{i})|=1$ for $i=1,2,3$. Let $S\cap N(x_{i})=\{z_{i}\}$. Then, $S=\{z_{1},z_{2},z_{3}\}$. So, $z_{i}\neq z_{j}$ for $i,j\in\{1,2,3\}$ and $i\neq j$ (else $|S|\leq 2$). Note that $S\cap N(x_{1})=\{z_{1}\}$. So, $x_{1}z_{2},x_{1}z_{3}\notin E(G)$. Using similar arguments, we can show that $x_{i}z_{j}\notin E(G)$ for $i,j\in\{1,2,3\}$ and $i\neq j$. Further, since $x_{2}x_{3}\in E(G)$ and $x_{3}z_{1}\notin E(G)$, $z_{1}\neq x_{2}$. Similarly, $z_{3}\neq x_{2}$. Also, note that if $z_{1}z_{3}\in E(G)$, then $z_{1}z_{2},z_{2}z_{3}\notin E(G)$ since $G[S]$ is $\{P_{3},K_{3}\}$-free by Observation 3. This together with the fact that $z_{3}x_{3}\in E(G)$ implies that $z_{2}\neq x_{3}$. Then, $\{z_{1},z_{3},x_{3}\}\cup\{z_{2}\}$ induces a ($P_{3}\cup K_{1}$) in $G$, a contradiction. So, $z_{1}z_{3}\notin E(G)$. But, then $\{z_{1},x_{1},x_{2}\}\cup\{z_{3}\}$ induces a ($P_{3}\cup K_{1}$), a contradiction and so $\rho^{o}(G)\leq 2$. Next, for every $r\geq 2$, we show that $\rho^{o}(G)\leq 2r$ for every connected $(P_{3}\cup rK_{1})$-free graph $G$.
@@
Let $G$ be a connected $(P_{3}\cup rK_{1})$-free graph. If $G$ does not contain any vertex subset that induces a $(P_{3}\cup(r-1)K_{1})$, then $G$ is $(P_{3}\cup(r-1)K_{1})$-free and so by induction assumption $\rho^{o}(G)\leq 2(r-1)<2r$. So, assume that $G$ has a vertex subset $D=\{x_{1},x_{2},x_{3},y_{1},y_{2},\ldots,y_{r-1}\}$ that induces a $(P_{3}\cup(r-1)K_{1})$ in $G$ with $x_{i}x_{i+1}\in E(G)$ for $i=1,2$. Since $G$ is $(P_{3}\cup rK_{1})$-free, every vertex in $V(G)\setminus D$ is adjacent to some vertex in $D$, i.e., $D$ is a dominating set in $G$. So, $V(G)=N[x_{1}]\cup N[x_{2}]\cup N[x_{3}]\cup N[y_{1}]\cup N[y_{2}]\cup\ldots\cup N[y_{r-1}]$. So, for any open packing $S$ in $G$, $S=\bigcup\limits_{u\in D}(S\cap N[u])$. Note that $|S\cap N[u]|\leq|S\cap N(u)|+|S\cap\{u\}|\leq 2$. Also, since $x_{1},x_{2},x_{3}\in N(x_{1})\cup N(x_{2})$, $\bigcup\limits_{i=1}^{3}N[x_{i}]=\bigcup\limits_{i=1}^{3}N(x_{i})$. So, $|S\cap(\bigcup\limits_{i=1}^{3}N[x_{i}])|=|S\cap(\bigcup\limits_{i=1}^{3}N(x_{i}))|\leq 3$. Therefore, $|S|=|S\cap V(G)|=|S\cap(N[x_{1}]\cup N[x_{2}]\cup N[x_{3}]\cup N[y_{1}]\cup N[y_{2}]\cup\ldots\cup N[y_{r-1}])|\leq 3+2(r-1)=2r+1$. Thus, to prove the claim, it is enough to contradict the possibility of $|S|=2r+1$. So, assume that there exists an open packing $S$ of size $12r+$ 1 in $G$. Then, by the above arguments, we can conclude that $|S\cap N[y_{j}]|=2$ for $j=1,2\ldots,r-1$ and $|S\cap(\bigcup\limits_{i=1}^{3}N(x_{i}))|=3$. Since $|S\cap N(y_{j})|\leq 1$ and $|S\cap N[y_{j}]|=2$, $y_{j}\in S$ and $S\cap N(y_{j})\neq\emptyset$. Let $S\cap N[y_{j}]=\{w_{j},y_{j}\}$ for $j=1,2,\ldots,r-1$ and let $S\cap(\bigcup\limits_{i=1}^{3}N(x_{i}))=\{z_{1},z_{2},z_{3}\}$ with $x_{i}z_{i}\in E(G)$ (then, since $|S\cap N(x_{i})|\leq 1$, we have $S\cap N(x_{i})=\{z_{i}\}$ for every $i=1,2,3$). So, $S=\left(\bigcup\limits_{i=1}^{3}\{z_{i}\}\right)\cup\left(\bigcup\limits_{j=1}^{r-1}\{y_{j}\}\right)\cup\left(\bigcup\limits_{j=1}^{r-1}\{w_{j}\}\right)$. This together with the fact that $|S|=2r+1$ imply that every vertex in the $(2r+1)$-tuple $(z_{1},z_{2},z_{3},y_{1},y_{2},\ldots,y_{r-1},w_{1},w_{2},\ldots,w_{r-1})$ is distinct. Further, since (i) $S\cap N(x_{i})=\{z_{i}\}$ and (ii) $S\cap N(y_{i})=\{w_{j}\}$, we have $z_{s}u\notin E(G)$ for $s\in\{1,2,3\}$ and $u\in D\setminus\{x_{s}\}$. We complete our claim in two cases based on $z_{1}z_{3}\in E(G)$ or not.
@@
Since $G[S]$ is $\{P_{3},K_{3}\}$-free by Observation 3, $z_{1}z_{2},z_{2}z_{3}\notin E(G)$. This implies that $z_{2}\neq x_{3}$. So, $\{x_{1},z_{1},z_{3}\}\cup\{z_{2},y_{1},y_{2},\ldots,y_{r-1}\}$ induces a $(P_{3}\cup rK_{1})$ in $G$, a contradiction.
@@
Therefore, $\rho^{o}(G)\leq 2r$. ∎
@@ Appendix C Proof Related to Section 4
Suppose that $S\subseteq V(G)$ is an independent set in $G$. On the contrary to Claim 7, assume that $S$ is not an open packing in $G^{\prime}$. Then, there exist vertices $u,u^{\prime}\in S$ and $w\in V(G^{\prime})$ such that $w\in N_{G^{\prime}}(u)\cap N_{G^{\prime}}(u^{\prime})$. Since $u,u^{\prime}\in V(G)\subseteq I$, $w\in C=E(G)\cup\{y,z\}$. By the construction of $G^{\prime}$, $uy,uz\notin E(G^{\prime})$ for any $u\in V(G)$. Thus, $w=e$ for some $e\in E(G)$. Then, by the construction of $G^{\prime}$, $e=uu^{\prime}\in E(G)$, a contradiction to $S$ being an independent set.
@@
Max-Open Packing is hard to approximate within a factor of $N^{(\frac{1}{2}-\epsilon)}$ for any $\epsilon>0$ in $K_{1,4}$-free split graphs unless $NP=P$, where $N$ is the number of vertices in $K_{1,4}$-free split graphs.
@@
$2(|S|-1)$ $>\frac{\rho^{o}(G^{\prime})}{(n^{2})^{\left(\frac{1}{2}-\epsilon\right)}}\hskip 14.22636pt\text{for }|S|\geq 2$
@@
The above inequality implies that Max-Independent Set has a $(n^{1-\epsilon})$-factor approximation algorithm which contradicts Theorem 4. ∎

### v2

@@ Appendix A Open Packing
Recall that $V(G_{r})=(\cup_{i=1}^{r}\{x_{i},y_{i},z_{i}\})\cup\{u,v\}$ and $E(G_{r})=(\cup_{i=1}^{r}\{x_{i}y_{i},y_{i}z_{i},z_{i}u\})\cup\{uv\}$. On the contrary to claim, assume that there exists a $(P_{4}\cup rK_{1})$, say $H$, in $G_{r}$. Let $\{a,b,c,d\}\in V(H)$ such that $G_{r}[{a,b,c,d}]\cong P_{4}$. Note that every $P_{4}$ in $G_{r}$ consists of the vertex $u$. So, $u\in\{a,b,c,d\}$. Similarly, observe that every $P_{4}$ in $G$ consists of exactly one vertex labelled $y_{i}$ for some $i=1,2,\ldots,r$. So, $y_{i}\in\{a,b,c,d\}$ for some $i=1,2,\ldots,r$. W.L.O.G., assume that $i=1$. Then, $\{a,b,c,d\}\in\{\{v,u,z_{1},y_{1}\},\{z_{j},u,z_{1},y_{1}\},\{u,z_{1},y_{1},x_{1}\}\}$. Since $H\cong(P_{4}\cup rK_{1})$ and $\{a,b,c,d\}\subseteq V(H)$ induces the $P_{4}$, $G_{r}[V(H)\setminus\{a,b,c,d\}]\cong rK_{1}$ and no vertex in $(V(H)\setminus\{a,b,c,d\})$ is adjacent to a vertex in $\{a,b,c,d\}$. So, for every possibility of $\{a,b,c,d\}$, $(V(H)\setminus\{a,b,c,d\})\subseteq\{x_{2},y_{2},\ldots,x_{r},y_{r}\}$. But the independence number of $G_{r}[\{x_{2},y_{2},\ldots,x_{r},y_{r}\}]$ is $(r-1)$, a contradiction to $(V(H)\setminus\{a,b,c,d\})\subseteq\{x_{2},y_{2},\ldots,x_{r},y_{r}\}$ and $(V(H)\setminus\{a,b,c,d\})\cong rK_{1}$, i.e., the independence number of an induced subgraph of $G_{r}[\{x_{2},y_{2},\ldots,x_{r},y_{r}\}]$ is at least $r$.
@@
We use induction on $r$ to prove this lemma as well. We begin with the proof that $\rho^{o}(G)\leq 2$ for every connected $(P_{3}\cup K_{1})$-free graph $G$. Note that, by Lemma 9, $\rho^{o}(G)\leq\gamma_{t}(G)\leq 3$, for every connected $(P_{3}\cup K_{1})$-free graph $G$. So, to prove the claim, it is enough to disprove the case $\rho^{o}(G)=3$. On the contrary, assume that there exists a $(P_{3}\cup K_{1})$-free graph $G$ with $\rho^{o}(G)=3$. Since the open packing number of a connected $P_{3}$-free graph is at most two by Remark 4 and $\rho^{o}(G)=3$, $G$ must contain $P_{3}$ as an induced subgraph. Let $D=\{x_{1},x_{2},x_{3}\}$ be a subset of $V(G)$ that induces a $P_{3}$ in $G$ with $x_{1}x_{2},x_{2}x_{3}\in E(G)$. Then, by the arguments in the proof of Lemma 9, $D$ is a total dominating set in $G$, i.e., $V(G)=\bigcup\limits_{i=1}^{3}N(x_{i})$. Then, for every open packing $S$ in $G$, $S=\bigcup\limits_{i=1}^{3}(S\cap N(x_{i}))$. We know that $|S\cap N(x_{i})|\leq 1$ for $i=1,2,3$. This, together with the fact that $S=\bigcup\limits_{i=1}^{3}(S\cap N(x_{i}))$, implies that for an open packing $S$ in $G$ of cardinality three, $|S\cap N(x_{i})|=1$ for $i=1,2,3$. Let $S\cap N(x_{i})=\{z_{i}\}$. Then, $S=\{z_{1},z_{2},z_{3}\}$. So, $z_{i}\neq z_{j}$ for $i,j\in\{1,2,3\}$ and $i\neq j$ (else $|S|\leq 2$). Note that $S\cap N(x_{1})=\{z_{1}\}$. So, $x_{1}z_{2},x_{1}z_{3}\notin E(G)$. Using similar arguments, we can show that $x_{i}z_{j}\notin E(G)$ for $i,j\in\{1,2,3\}$ and $i\neq j$. Further, since $x_{2}x_{3}\in E(G)$ and $x_{3}z_{1}\notin E(G)$, $z_{1}\neq x_{2}$. Similarly, $z_{3}\neq x_{2}$. Also, note that if $z_{1}z_{3}\in E(G)$, then $z_{1}z_{2},z_{2}z_{3}\notin E(G)$ since $G[S]$ is $\{P_{3},K_{3}\}$-free by Observation 3. This, together with the fact that $z_{3}x_{3}\in E(G)$, implies that $z_{2}\neq x_{3}$. Then, $\{z_{1},z_{3},x_{3}\}\cup\{z_{2}\}$ induces a ($P_{3}\cup K_{1}$) in $G$, a contradiction. So, $z_{1}z_{3}\notin E(G)$. But, then $\{z_{1},x_{1},x_{2}\}\cup\{z_{3}\}$ induces a ($P_{3}\cup K_{1}$), a contradiction. So, $\rho^{o}(G)\leq 2$. Next, for every $r\geq 2$, we show that $\rho^{o}(G)\leq 2r$ for every connected $(P_{3}\cup rK_{1})$-free graph $G$.
@@
Let $G$ be a connected $(P_{3}\cup rK_{1})$-free graph. If $G$ does not contain any vertex subset that induces a $(P_{3}\cup(r-1)K_{1})$, then $G$ is $(P_{3}\cup(r-1)K_{1})$-free, and so by induction assumption $\rho^{o}(G)\leq 2(r-1)<2r$. So, assume that $G$ has a vertex subset $D=\{x_{1},x_{2},x_{3},y_{1},y_{2},\ldots,y_{r-1}\}$ that induces a $(P_{3}\cup(r-1)K_{1})$ in $G$ with $x_{i}x_{i+1}\in E(G)$ for $i=1,2$. Since $G$ is $(P_{3}\cup rK_{1})$-free, every vertex in $V(G)\setminus D$ is adjacent to some vertex in $D$, i.e., $D$ is a dominating set in $G$. So, $V(G)=N[x_{1}]\cup N[x_{2}]\cup N[x_{3}]\cup N[y_{1}]\cup N[y_{2}]\cup\ldots\cup N[y_{r-1}]$. This implies that, for any open packing $S$ in $G$, $S=\bigcup\limits_{u\in D}(S\cap N[u])$. Note that $|S\cap N[u]|\leq|S\cap N(u)|+|S\cap\{u\}|\leq 2$. Also, since $x_{1},x_{2},x_{3}\in N(x_{1})\cup N(x_{2})$, $\bigcup\limits_{i=1}^{3}N[x_{i}]=\bigcup\limits_{i=1}^{3}N(x_{i})$. So, $|S\cap(\bigcup\limits_{i=1}^{3}N[x_{i}])|=|S\cap(\bigcup\limits_{i=1}^{3}N(x_{i}))|\leq 3$. Therefore, $|S|=|S\cap V(G)|=|S\cap(N[x_{1}]\cup N[x_{2}]\cup N[x_{3}]\cup N[y_{1}]\cup N[y_{2}]\cup\ldots\cup N[y_{r-1}])|\leq 3+2(r-1)=2r+1$. Thus, to prove the claim, it is enough to contradict the possibility of $|S|=2r+1$. Assume that there exists an open packing $S$ of size $12r+$ 1 in $G$. Then, by the above arguments, we can conclude that $|S\cap N[y_{j}]|=2$ for $j=1,2\ldots,r-1$ and $|S\cap(\bigcup\limits_{i=1}^{3}N(x_{i}))|=3$. Since $|S\cap N(y_{j})|\leq 1$ and $|S\cap N[y_{j}]|=2$, $y_{j}\in S$ and $S\cap N(y_{j})\neq\emptyset$. Let $S\cap N[y_{j}]=\{w_{j},y_{j}\}$ for $j=1,2,\ldots,r-1$ and let $S\cap(\bigcup\limits_{i=1}^{3}N(x_{i}))=\{z_{1},z_{2},z_{3}\}$ with $x_{i}z_{i}\in E(G)$ (then, since $|S\cap N(x_{i})|\leq 1$, we have $S\cap N(x_{i})=\{z_{i}\}$ for every $i=1,2,3$). So, $S=\left(\bigcup\limits_{i=1}^{3}\{z_{i}\}\right)\cup\left(\bigcup\limits_{j=1}^{r-1}\{y_{j}\}\right)\cup\left(\bigcup\limits_{j=1}^{r-1}\{w_{j}\}\right)$. This, together with the fact that $|S|=2r+1$, implies that every vertex in the $(2r+1)$-tuple $(z_{1},z_{2},z_{3},y_{1},y_{2},\ldots,y_{r-1},w_{1},w_{2},\ldots,w_{r-1})$ is distinct. Further, since (i) $S\cap N(x_{i})=\{z_{i}\}$ and (ii) $S\cap N(y_{i})=\{w_{j}\}$, we have $z_{s}u\notin E(G)$ for $s\in\{1,2,3\}$ and $u\in D\setminus\{x_{s}\}$. We complete our claim in two cases based on $z_{1}z_{3}\in E(G)$ or not.
@@
Then, by Observation 3, $z_{1}z_{2},z_{2}z_{3}\notin E(G)$. This implies that $z_{2}\neq x_{3}$. So, $\{x_{1},z_{1},z_{3}\}\cup\{z_{2},y_{1},y_{2},\ldots,y_{r-1}\}$ induces a $(P_{3}\cup rK_{1})$ in $G$, a contradiction.
@@
In both cases, we arrive at a contradiction. Therefore, $\rho^{o}(G)\leq 2r$. ∎
@@ Appendix C Proof Related to Section 4
Suppose that $S\subseteq V(G)$ is an independent set in $G$. On the contrary to Claim 7, assume that $S$ is not an open packing in $G^{\prime}$. Then, there exist vertices $u,u^{\prime}\in S$ and $w\in V(G^{\prime})$ such that $w\in N_{G^{\prime}}(u)\cap N_{G^{\prime}}(u^{\prime})$. Since $u,u^{\prime}\in V(G)\subseteq I$, $w\in C=E(G)\cup\{y,z\}$. By the construction of $G^{\prime}$, $uy,uz\notin E(G^{\prime})$ for any $u\in V(G)$. Thus, $w=e$ for some $e\in E(G)$. Then, by the construction of $G^{\prime}$, $e=uu^{\prime}\in E(G)$, a contradiction to $S$ being an independent set in $G$.
@@
Max-Open Packing is hard to approximate within a factor of $N^{(\frac{1}{2}-\epsilon)}$ for any $\epsilon>0$ in $K_{1,4}$-free split graphs unless P = NP, where $N$ is the number of vertices in $K_{1,4}$-free split graphs.
@@
$2(|S|-1)$ $>\frac{\rho^{o}(G^{\prime})}{(n^{2})^{\left(\frac{1}{2}-\epsilon\right)}}\hskip 14.22636pt\text{for }|S|\geq 2\text{ and }n\geq 4$
@@
The above inequality implies that Max-Independent Set has a $(n^{1-\epsilon})$-factor approximation algorithm, which contradicts Theorem 4. ∎